# Christian Shelton, Manmohan Chandraker‡,⋆, Abhishek Aich

### v1

@@ 1 Introduction
Grounding large language models (LLMs) to domain-specific requirements remains a significant challenge due to their general-purpose training and lack of inductive bias toward structured domain knowledge [1]. LLMs are often pretrained on broad internet data, which may lead to imprecise outputs when applied to specialized domains such as driving video understanding [2]. Furthermore, the lack of interpretability and limited mechanisms for precise reasoning make it difficult to guarantee reliability in high-stakes applications [3]. In this paper, we tackle the problem of generating causally and spatially grounded LLM responses to user queries, given a front-view dash-cam video. The objective is to produce explanations that accurately reflect the underlying events in the scene.
The proliferation of advanced driver assistance systems has led to an abundance of video data, often serving as the primary or sole modality for system analysis and validation. In many consumer-grade and fleet-level applications, supplementary sensors like LiDAR, GPS, or CAN bus data are either unavailable or impractical to collect due to cost and integration complexities [4, 5]. Consequently, extracting meaningful insights from visual spatio-temporal data becomes crucial for assessing system responses, diagnosing failures, and refining vehicle performance [6, 7, 8, 9, 10, 11]. However, developing video analysis systems that generate grounded and reliable responses solely from camera input remains a formidable challenge, without external supervision [12, 13].
A seemingly straightforward solution to address the above problem is to develop end-to-end video-based Vision-Language Models (V-VLMs) [14, 15, 16]. However, although existing V-VLMs generate reasonable responses, they exhibit limitations in spatial reasoning, causal inference, and fine-grained scene understanding [17, 18, 19] as shown in Figure 1. This means misinterpretation of critical visual cues can lead to incorrect conclusions about hazards, traffic signals, or object presence. Furthermore, incorporating new functionality without requiring extensive retraining or fine-tuning is difficult with VLMs [20].
@@
Figure 1: Advantages of $\bm{\dot{\imath}}$Finder. Baselines VideoLLaMA2 [14], VideoLLaVA [15], and DriveMM [16] struggle with spatial reasoning, and fine-grained scene understanding, misinterpreting critical cues. $\bm{\dot{\imath}}$Finder’s structured scene approach mitigates these errors for more accurate responses.
To address these limitations, we build on the principle that perception should be decoupled from LLM reasoning. In particular, we propose $\bm{\dot{\imath}}$Finder, a vision-language pipeline that extracts driving domain-relevant visual cues and passes them to the LLM via structured prompts, enabling post-hoc scenario understanding through symbolic, temporally grounded reasoning. This structured scene representations encode dynamic object pose, orientation, and semantic lane context in a hierarchical format, enabling symbolic reasoning over frame-indexed data. $\bm{\dot{\imath}}$Finder leverages pre-trained vision models to extract domain-relevant cues and organizes them into a hierarchical data format. This structured input enables the LLM to reason accurately about driving scenarios, correcting or augmenting generic V-VLM outputs with grounded, verifiable evidence tailored to the driving video context. For example, in Figure˜1, we can observe that unlike baselines [16, 21, 14, 15] that provide incorrect or generic explanations, $\bm{\dot{\imath}}$Finder correctly identifies the white car’s lane-cutting maneuver, aligning with the ground truth. It leverages the data structure of the input video and uses object tracking (‘Object ID 13’), distance change (‘9.78m $\to$ 6.14m’), and orientation (‘rot_y = -2.1 to indicate a left turn’) to draw conclusions. Our approach not only mitigates the limitations of end-to-end V-VLMs but also improves reliability and transparency.
@@
  • Improved Performance via Structured Grounding. By decoupled perception and LLM reasoning, and grounding LLMs in hierarchical, interpretable video representations, $\bm{\dot{\imath}}$Finder beats both generalist and driving-specific VLMs, without any fine-tuning. $\bm{\dot{\imath}}$Finder, for example, in the accident reasoning dataset MM-AU [22], beats the best performing generalist V-VLM by 10.5% and driving-specialized V-VLM by 39.17%.
  • Enhanced Explainability through Explicit Reasoning. Unlike end-to-end VLMs that rely on implicit cues, $\bm{\dot{\imath}}$Finder provides symbolic cues—such as object orientation, lane context, and distance—that allows the LLM to generate transparent, verifiable explanations, as shown in Figure˜3.
  • Object-orientation and Global Environmental Dominate. Surprisingly, our ablation studies in Table˜7 reveal that object-orientation and global environmental context contribute more to reliable post-hoc reasoning than other factors like distance and lane location. For instance, removing object orientation information led to a significant drop in reasoning accuracy ($\sim$4.5%), while removing distance or lane location had a comparatively minor effect ($\sim$2.7%).
@@ 2 Related Works
Post-hoc driving video analysis for front-cam setting (rather than surveillance-cam setting, such as in traffic intersections [23]) has recently become a strong research interest. Prior works on driving video analysis [24, 22] have been dominantly focused on accidents and their possible prevention analysis. Different from these, [25] provides a detailed benchmark in which models can be tested on their understanding of the ego-vehicle’s perspective based on their understanding of dynamic scenes, rather than merely their ability to describe the video event. In this work and unlike prior driving VLMs [18, 16], we develop a vision-language pipeline that solely focuses on such offline driving video analysis.
@@
The success of the image-bases VLSs has sparked a growing interest in the video-based VLMs within the research community. In the video-language domain, the prevailing approach now involves large-scale video-text pre-training followed by fine-tuning for specific tasks [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 21, 41, 42, 43, 44, 45, 15]. Within the driving video domain, multiple image-language models have been introduced [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] for various driving-related tasks. Although these methods perform well, their effectiveness is restricted to specific scenarios (Bird’s-Eye-View-based representations, multi-view representations, etc.) and driving-specific objectives that dominantly do not aid offline video analysis tasks. For video-domain driving specific oundational models, methods like [16, 57, 18] have been introduced, however, they lack the capability to process different driving video with fine-grained analysis like ego-car attributes. Moreover, these VLMs are designed for tasks to represent in-moment vehicle driver, rather than video analysis.
@@ 3.1 Proposed Framework
Finally, the structured data $\mathcal{D}$} is passed to an LLM, denoted ${\mathcal{F}}_{\text{LLM}}$, along with a prompt $P_{\text{LLM}}$, to generate the final response $\bm{R}$} to the user’s query.
Next, we describe the function $\mathcal{F}$} that aims to extract $\bm{D}$} in an LLM-friendly manner. In practice, $\mathcal{F}$} is decomposed into multiple specialized modules: 2D/3D object detection, lane localization, distance estimation, attribute estimation, ego-state estimation and scene understanding. Each of the modules is instantiated with popular pre-trained open-source models, requiring no fine-tuning when used within $\bm{\dot{\imath}}$Finder’s pipeline.
@@ 3.1.1 Visual Information Extraction
Front-view dash-cam videos often exhibit distortions due to the characteristics of front-facing cameras [58]. To address this and ensure optimal performance for perception models to follow, $\bm{\dot{\imath}}$Finder begins by correcting these distortions through the estimation of camera intrinsics $\bm{K}$} (focal length and principal point), and distortion coefficients, including radial distortion $k_{1},k_{2},k_{3}]$ and tangential distortion $p_{1},p_{2}]$ coefficients. To correct for lens distortion, we compute a mapping function $\mathcal{R}$} that transforms coordinates in the rectified (undistorted) image $\bm{I}$} to corresponding coordinates in the distorted image ${\bm{I}}_{d}$. The rectified image is obtained by:
@@
    Next, $\Delta\theta_{t}$ is used to classify the vehicle’s turn into the three categories ($\tau_{a}$ as threshold) as
@@
This step estimates the distance of the objects in the scene w.r.t. ego-vehicle from the video.
@@
See Supplementary Material for a qualitative illustration of Step 5 and Step 6.
@@
  • Intuition. Object attributes enhance perception with human-like reasoning, useful for scene interpretation, and understanding high-level decision-making.
@@
  • Intuition. From Step 4-8, all extracted information was from a 2D perspective. However, a critical aspect of the objects missing are their orientation as per the ego-vehicle’s view.
  • Method. Using a 3D detection model ${\mathcal{F}}_{\text{3D-det}}$, we predict $p_{t}$ 3D bounding boxes and extract the yaw $\theta_{t,i}\in[-\pi,\pi]$ for each object. as ${\mathcal{F}}_{\text{3D-det}}:({\bm{I}}_{t})\to\{\theta_{t,i}\}_{i=1}^{p_{t}}$. Next, we project each 3D bounding box into 2D image space using the camera intrinsic matrix $\bm{K}$} from Step 1. We then apply the Hungarian algorithm [59] to match projected boxes with detected objects and transfer $\theta_{t,i}$ to the corresponding local object.
@@ 3.1.2 Proposed Data Structure and Prompt
The prompt $P_{\text{LLM}}$ is designed as a three-block structure to optimize the model’s grounded reasoning. The first component, Key Explanation, provides a precise and explicit interpretation of the scene representation $\mathcal{D}$}, reducing ambiguity in the model’s input. Prior work shows that LLMs benefit from well-disambiguated inputs when handling symbolic data [60, 61]. The second component, Step Instructions, decomposes the reasoning task into explicit sub-goals. This design is motivated by findings in cognitive science and neural model alignment, where step-by-step prompting improves reasoning accuracy and consistency [62, 63]. The third component, Peer Instruction, informs the model that peer-generated answers may be unreliable and explicitly encourages independent reasoning. Together, these three components operationalize input grounding, procedural reasoning, and epistemic caution—three necessary conditions for robust and generalizable performance in tasks involving multi-step inference from structured dynamic scenes. Our prompt has been provided in the Supplementary Material.
@@ 4 Experiments
$\bm{\dot{\imath}}$Finder leverages a combination of state-of-the-art models in each of its steps. In Step 1, we use GeoCalib [64] for estimating camera parameters and distortion coefficients. To correct the lens distortion, we use OpenCV’s undistort [65] function for $\mathcal{R}$}. In Step 2, we use InternVL [66] for ${\mathcal{F}}_{\text{I-VLM}}$ and VideoLLaMA2 [14] for ${\mathcal{F}}_{\text{V-VLM}}$. In Step 3, we use DROID-SLAM [67] for ${\mathcal{F}}_{\text{cam-pose}}$. In Step 4 for ${\mathcal{F}}_{\text{2D-det}}$, we use OWL-V2 [68] and ByteTracker [69] for ${\mathcal{F}}_{\text{2D-track}}$. In Step 5, we use OMR [70] for ${\mathcal{F}}_{\text{lane}}$. In Step 6, Metric3D [71] is used for ${\mathcal{F}}_{\text{depth}}$ and SAM [72] for ${\mathcal{F}}_{\text{seg}}$. In Step 7, we again use InternVL for ${\mathcal{F}}_{\text{I-VLM}}$. In Step 8, we use CenterTrack [73] for ${\mathcal{F}}_{\text{3D-det}}$. For peer-informed reasoning, VideoLLaMA2 [14] serves as the default peer model unless otherwise specified. For final reasoning step, we use GPT-4o-mini [74] for ${\mathcal{F}}_{\text{LLM}}$. The full list of hyperparameters and prompts is in the Supplementary Material. Note that our goal is not to compare model variants, but to demonstrate the effectiveness of the combined vision–language pipeline. Each can be easily swapped for stronger versions for further improvements.
@@
We choose datasets and baselines that require all the methods to analyze the complete video before answering the user query. To this end, we use four benchmarks: MM-AU (Multi-Modal Accident Video Understanding) [22], SUTD (Traffic Question Answering) [75], LingoQA [76], and Nexar dataset [77]. For baselines, we compare $\bm{\dot{\imath}}$Finder with V-VLMs that are trained to provide open-ended responses to users’ queries on input videos. In particular, we compare against recent state-of-the-art general V-VLMs VideoLLaMA2 [14], VideoChat2 [21], and VideoLLaVA [15], as well as those proposed for driving video analysis DriveMM [16] and WiseAD [18]. The dataset and evaluation metric details are provided in the Supplementary Material. All the following evaluations are in a strictly zero-shot setting, with no fine-tuning.
@@
We analyze the performance on accident-cause and traffic scene understanding on MM-AU and SUTD datasets under multi-choice VQA setup, shown in Table˜3 and Table˜5, and gain three insights on $\bm{\dot{\imath}}$Finder versatility and performance. One, $\bm{\dot{\imath}}$Finder outperforms both general-purpose and driving-specialized models on MM-AU and SUTD without any fine-tuning, highlighting its strong out-of-the-box reasoning capabilities. Two, it achieves the best accuracy in all six SUTD categories, from basic understanding to attribution, showing robust generalization across diverse cognitive tasks. Three, the 39-point gap between $\bm{\dot{\imath}}$Finder and DriveMM on MM-AU underscores the limitations of current domain-specific models and the strength of $\bm{\dot{\imath}}$Finder’s architecture in complex, real-world driving scenarios.
@@
VideoChat2 [21] 49.56 42.17
Video-LLaVA [15] 43.63 38.35
@@
VideoLLaMA2 [14] 36.00 4.15 33.45 26.28
VideoChat2 [21] 41.20 6.58 36.81 40.98
Video-LLaVA [15] 21.00 4.26 26.99 31.23
@@
VideoChat2 [21] 58.0 0.62 0.57 0.68
VideoLLaMA2 [14] 50.0 0.67 0.50 1.00
@@
DriveMM [16] 49.0 0.32 0.48 0.24
WiseAD [18] 50.0 0.67 0.50 1.00
@@
Method U F R C I A
Generalist Models
VideoLLaMA2 [14] 49.2 39.0 48.5 53.5 35.8 45.2
VideoChat2 [21] 42.5 38.1 43.8 49.2 30.4 42.8
Video-LLaVA [15] 39.7 37.2 35.8 40.5 31.1 36.4
Driving-specialized Methods
DriveMM [16] 47.6 38.6 40.1 43.2 38.5 37.7
$\bm{\dot{\imath}}$Finder (Ours) 52.2 43.5 50.2 56.8 39.2 49.6
Table 5: Result breakdown on MMAU weather categories. $\bm{\dot{\imath}}$Finder achieves best performance across adversarial conditions showing its adaptability to diverse environments.
Method Foggy Rainy Snowy Sunny Day Night
Generalist Models
VideoLLaMA2 (w/pt) [14] 66.67 44.83 40.25 54.56 53.23 50.23
VideoChat2 [21] 58.33 54.31 49.69 49.16 50.29 43.84
Video-LLaVA [15] 58.33 41.38 42.77 43.76 43.25 46.58
Driving-specialized Methods
DriveMM [16] 33.33 24.14 20.13 24.55 23.41 30.59
$\bm{\dot{\imath}}$Finder (Ours) 75.00 65.52 57.86 63.69 63.32 63.93
Next, we analyzed all methods under an open-ended VQA setup on the LingoQA dataset in Table˜3 and provided the following insights. One, $\bm{\dot{\imath}}$Finder achieves the best Lingo-Judge accuracy among all evaluated methods, outperforming VideoLLaMA2 by 3%, VideoChat2 by 8%, and Video-LLaVA by 23.2%. Further, it outperforms driving-specialized models, WiseAD by 30.8%. Two, while our method performs competitively on BLEU and METEOR metrics, it slightly lags behind VideoChat2. This suggests that VideoChat2 generates responses that are more lexically similar to reference answers but do not necessarily reflect greater factual correctness. We also show this in Figure˜3.
Finally, we analyze all methods for accident occurrence prediction on the Nexar dataset in Table˜3. $\bm{\dot{\imath}}$Finder achieves the highest accuracy at 62.0%, outperforming both generalist models, as well as driving-specialized models. Notably, while VideoLLaMA2 and WiseAD achieve high F1 and perfect recall by always predicting accidents, this limits real-world reliability. In contrast, VideoChat2 balances precision and recall for better accuracy. Building on this, $\bm{\dot{\imath}}$Finder integrates it within the peer V-VLM to enhance both accuracy and precision.

### v3

@@ 1 Introduction
Grounding large language models (LLMs) to domain-specific requirements remains a significant challenge due to their general-purpose training and lack of inductive bias toward structured domain knowledge [bommasani2021opportunities]. LLMs are often pretrained on broad internet data, which may lead to imprecise outputs when applied to specialized domains such as driving video understanding [nori2023capabilities]. Furthermore, the lack of interpretability and limited mechanisms for precise reasoning make it difficult to guarantee reliability in high-stakes applications [binz2023robustness]. In this paper, we tackle the problem of generating causally and spatially grounded LLM responses to user queries, given a front-view dash-cam video. The objective is to produce explanations that accurately reflect the underlying events in the scene.
The proliferation of advanced driver assistance systems has led to an abundance of video data, often serving as the primary or sole modality for system analysis and validation. In many consumer-grade and fleet-level applications, supplementary sensors like LiDAR, GPS, or CAN bus data are either unavailable or impractical to collect due to cost and integration complexities [aptiv2023transition, andres2017data]. Consequently, extracting meaningful insights from visual spatio-temporal data becomes crucial for assessing system responses, diagnosing failures, and refining vehicle performance [poms2018scanner, bojarski2016end, chen2022milestones, baran2022prediction, lin2024harnessing, zhou2024gpt]. However, developing video analysis systems that generate grounded and reliable responses solely from camera input remains a formidable challenge, without external supervision [fu2024drive, fernando2022automated].
A seemingly straightforward solution to address the above problem is to develop end-to-end video-based Vision-Language Models (V-VLMs) [damonlpsg2024videollama2, lin2023video, huang2024drivemm]. However, although existing V-VLMs generate reasonable responses, they exhibit limitations in spatial reasoning, causal inference, and fine-grained scene understanding [wen2023road, zhang2024wisead, liu2025vlme2eenhancingendtoendautonomous] as shown in Figure 1. This means misinterpretation of critical visual cues can lead to incorrect conclusions about hazards, traffic signals, or object presence. Furthermore, incorporating new functionality without requiring extensive retraining or fine-tuning is difficult with VLMs [xu2024vlmadendtoendautonomousdriving].
@@
Figure 1: Advantages of $\bm{\dot{\imath}}$Finder. Baselines VideoLLaMA2 [damonlpsg2024videollama2], VideoLLaVA [lin2023video], and DriveMM [huang2024drivemm] struggle with spatial reasoning, and fine-grained scene understanding, misinterpreting critical cues. $\bm{\dot{\imath}}$Finder’s structured scene approach mitigates these errors for more accurate responses.
To address these limitations, we build on the principle that perception should be decoupled from LLM reasoning. In particular, we propose $\bm{\dot{\imath}}$Finder, a vision-language pipeline that extracts driving domain-relevant visual cues and passes them to the LLM via structured prompts, enabling post-hoc scenario understanding through symbolic, temporally grounded reasoning. This structured scene representations encode dynamic object pose, orientation, and semantic lane context in a hierarchical format, enabling symbolic reasoning over frame-indexed data. $\bm{\dot{\imath}}$Finder leverages pre-trained vision models to extract domain-relevant cues and organizes them into a hierarchical data format. This structured input enables the LLM to reason accurately about driving scenarios, correcting or augmenting generic V-VLM outputs with grounded, verifiable evidence tailored to the driving video context. For example, in Figure˜1, we can observe that unlike baselines [huang2024drivemm, li2023videochat, damonlpsg2024videollama2, lin2023video] that provide incorrect or generic explanations, $\bm{\dot{\imath}}$Finder correctly identifies the white car’s lane-cutting maneuver, aligning with the ground truth. It leverages the data structure of the input video and uses object tracking (‘Object ID 13’), distance change (‘9.78m $\to$ 6.14m’), and orientation (‘rot_y = -2.1 to indicate a left turn’) to draw conclusions. Our approach not only mitigates the limitations of end-to-end V-VLMs but also improves reliability and transparency.
@@
  • Improved Performance via Structured Grounding. By decoupled perception and LLM reasoning, and grounding LLMs in hierarchical, interpretable video representations, $\bm{\dot{\imath}}$Finder beats both generalist and driving-specific VLMs, without any fine-tuning. $\bm{\dot{\imath}}$Finder, for example, in the accident reasoning dataset MM-AU [fang2024abductive], beats the best performing generalist V-VLM by 10.5% and driving-specialized V-VLM by 39.17%.
  • Enhanced Explainability through Explicit Reasoning. Unlike end-to-end VLMs that rely on implicit cues, $\bm{\dot{\imath}}$Finder provides symbolic cues—such as object orientation, lane context, and distance—that allows the LLM to generate transparent, verifiable explanations, as shown in Figure˜5.
  • Object-orientation and Global Environmental Dominate. Surprisingly, our ablation studies in Table˜6 reveal that object-orientation and global environmental context contribute more to reliable post-hoc reasoning than other factors like distance and lane location. For instance, removing object orientation information led to a significant drop in reasoning accuracy ($\sim$4.5%), while removing distance or lane location had a comparatively minor effect ($\sim$2.7%).
@@ 2 Related Works
Post-hoc driving video analysis for front-cam setting (rather than surveillance-cam setting, such as in traffic intersections [xu2024tad]) has recently become a strong research interest. Prior works on driving video analysis [kim2019crash, fang2024abductive] have been dominantly focused on accidents and their possible prevention analysis. Different from these, [singh2022road] provides a detailed benchmark in which models can be tested on their understanding of the ego-vehicle’s perspective based on their understanding of dynamic scenes, rather than merely their ability to describe the video event. Recent work has explored zero-shot hazard identification and out-of-distribution challenges, with [alshami2024coool] introducing the COOOL benchmark for evaluating autonomous driving models on out-of-label objects, [picek2025zero] demonstrating zero-shot hazard identification approaches on this benchmark, and [shriram2025towards] proposing multi-agent vision-language systems for detecting novel hazardous objects. In this work, unlike prior driving VLMs [zhang2024wisead, huang2024drivemm], we develop a vision-language pipeline that solely focuses on such offline driving video analysis.
@@
The success of the image-bases VLSs has sparked a growing interest in the video-based VLMs within the research community. In the video-language domain, the prevailing approach now involves large-scale video-text pre-training followed by fine-tuning for specific tasks [MiechASLSZ20, cpd, merlot, umt, uniformerv2, videoclip, hu2022scaling, dou2022empirical, shen2021much, yao2021filip, videobert, actbert, wang2022internvideo, chen2022internvideo, maaz2023video, li2023videochat, li2024mvbench, gao2023llama, li2024llama, zhang2023video, jin2024chat, lin2023video]. Within the driving video domain, multiple image-language models have been introduced [wen2023dilu, mao2023gpt, xu2024drivegpt4, huang2025making, shao2024lmdrive, wang2023drivemlm, nie2025reason2drive, ding2024holistic, wang2024omnidrive, zhou2024embodied, chen2024automated] for various driving-related tasks. Although these methods perform well, their effectiveness is restricted to specific scenarios (Bird’s-Eye-View-based representations, multi-view representations, etc.) and driving-specific objectives that dominantly do not aid offline video analysis tasks. For video-domain driving specific foundational models, methods like [huang2024drivemm, dinh2024trafficvlm, zhang2024wisead] have been introduced. However, they lack the capability to process different driving video with fine-grained analysis like ego-car attributes. Moreover, these VLMs are designed for tasks to represent in-moment vehicle driver, rather than video analysis.
@@ 3.1 Proposed Framework
Finally, the structured data $\mathcal{D}$} is passed to an LLM, denoted ${\mathcal{F}}_{\text{LLM}}$, along with a prompt $P_{\text{LLM}}$, to generate the final response $\bm{R}$} to the user’s query. Next, we describe the function $\mathcal{F}$} that aims to extract $\bm{D}$} in an LLM-friendly manner. In practice, $\mathcal{F}$} is decomposed into multiple specialized modules: 2D/3D object detection, lane localization, distance estimation, attribute estimation, ego-state estimation, and scene understanding. Each of the modules is instantiated with popular pre-trained open-source models, requiring no fine-tuning when used within $\bm{\dot{\imath}}$Finder’s pipeline.
@@ 3.1.1 Visual Information Extraction
Front-view dash-cam videos often exhibit distortions due to the characteristics of front-facing cameras [giovannini2021importance]. To address this and ensure optimal performance for perception models to follow, $\bm{\dot{\imath}}$Finder begins by correcting these distortions through the estimation of camera intrinsics $\bm{K}$} (focal length and principal point), and distortion coefficients, including radial distortion $k_{1},k_{2},k_{3}]$ and tangential distortion $p_{1},p_{2}]$ coefficients. To correct for lens distortion, we compute a mapping function $\mathcal{R}$} that transforms coordinates in the rectified (undistorted) image $\bm{I}$} to corresponding coordinates in the distorted image ${\bm{I}}_{d}$. The rectified image is obtained by:
@@
    Next, $\Delta\theta_{t}$ is used to classify the vehicle’s turn into the three categories ($\tau_{a}$ as a threshold) as
@@
Figure 3: Lane location estimation. Detected objects are assigned a lane by mapping the bottom midpoint of the corresponding bounding box (bottom middle point of image for ego) to sections identified by the lane detection model.
Figure 4: Distance estimation. Each object’s distance is determined by averaging the depth values within its segmented region.
@@
This step estimates the distance of the objects in the scene w.r.t. the ego-vehicle from the video.
@@
A qualitative illustration of Step 5 and Step 6 is shown in Figure˜4 and Figure˜4, respectively.
@@
  • Intuition. Object attributes enhance perception with human-like reasoning, useful for scene interpretation and understanding high-level decision-making.
@@
  • Intuition. From Step 4-8, all extracted information was from a 2D perspective. However, a critical aspect of the objects missing is their orientation as per the ego-vehicle’s view.
  • Method. Using a 3D detection model ${\mathcal{F}}_{\text{3D-det}}$, we predict $p_{t}$ 3D bounding boxes and extract the yaw $\theta_{t,i}\in[-\pi,\pi]$ for each object. as ${\mathcal{F}}_{\text{3D-det}}:({\bm{I}}_{t})\to\{\theta_{t,i}\}_{i=1}^{p_{t}}$. Next, we project each 3D bounding box into 2D image space using the camera intrinsic matrix $\bm{K}$} from Step 1. We then apply the Hungarian algorithm [kuhn1955hungarian] to match projected boxes with detected objects and transfer $\theta_{t,i}$ to the corresponding local object.
@@ 3.1.2 Proposed Data Structure and Prompt
The prompt $P_{\text{LLM}}$ is designed as a three-block structure to optimize the model’s grounded reasoning. The first component, Key Explanation, provides a precise and explicit interpretation of the scene representation $\mathcal{D}$}, reducing ambiguity in the model’s input. Prior work shows that LLMs benefit from well-disambiguated inputs when handling symbolic data [zhou2023least, webson2022prompt]. The second component, Step Instructions, decomposes the reasoning task into explicit sub-goals. This design is motivated by findings in cognitive science and neural model alignment, where step-by-step prompting improves reasoning accuracy and consistency [wei2022chain, nye2021show]. The third component, Peer Instruction, informs the model that peer-generated answers may be unreliable and explicitly encourages independent reasoning. Together, these three components operationalize input grounding, procedural reasoning, and epistemic caution—three necessary conditions for robust and generalizable performance in tasks involving multi-step inference from structured dynamic scenes. Our prompt has been provided in the Supplementary Material.
@@ 4 Experiments
$\bm{\dot{\imath}}$Finder leverages a combination of state-of-the-art models in each of its steps. In Step 1, we use GeoCalib [veicht2024geocalib] for estimating camera parameters and distortion coefficients. To correct the lens distortion, we use OpenCV’s undistort [opencv_undistort] function for $\mathcal{R}$}. In Step 2, we use InternVL [chen2024internvl] for ${\mathcal{F}}_{\text{I-VLM}}$ and VideoLLaMA2 [damonlpsg2024videollama2] for ${\mathcal{F}}_{\text{V-VLM}}$. In Step 3, we use DROID-SLAM [teed2021droid] for ${\mathcal{F}}_{\text{cam-pose}}$. In Step 4 for ${\mathcal{F}}_{\text{2D-det}}$, we use OWL-V2 [minderer2024scalingopenvocabularyobjectdetection] and ByteTracker [zhang2022bytetrack] for ${\mathcal{F}}_{\text{2D-track}}$. In Step 5, we use OMR [Jin2024omr] for ${\mathcal{F}}_{\text{lane}}$. In Step 6, Metric3D [hu2024metric3dv2] is used for ${\mathcal{F}}_{\text{depth}}$ and SAM [kirillov2023segment] for ${\mathcal{F}}_{\text{seg}}$. In Step 7, we again use InternVL for ${\mathcal{F}}_{\text{I-VLM}}$. In Step 8, we use CenterTrack [zhou2020tracking] for ${\mathcal{F}}_{\text{3D-det}}$. For peer-informed reasoning, VideoLLaMA2 [damonlpsg2024videollama2] serves as the default peer model unless otherwise specified. For final reasoning step, we use GPT-4o-mini [openai2024gpt4ocard] for ${\mathcal{F}}_{\text{LLM}}$. The full list of hyperparameters and prompts is in the Supplementary Material. Note that our goal is not to compare model variants, but to demonstrate the effectiveness of the combined vision–language pipeline. Each can be easily swapped for stronger versions for further improvements.
@@
We choose datasets and baselines that require all the methods to analyze the complete video before answering the user query. To this end, we use four benchmarks: MM-AU (Multi-Modal Accident Video Understanding) [fang2024abductive], SUTD (Traffic Question Answering) [Xu_2021_CVPR], LingoQA [marcu2024lingoqa], and Nexar [nexar_dataset] dataset. For baselines, we compare $\bm{\dot{\imath}}$Finder with V-VLMs that are trained to provide open-ended responses to users’ queries on input videos. In particular, we compare against recent state-of-the-art general V-VLMs VideoLLaMA2 [damonlpsg2024videollama2], VideoChat2 [li2023videochat], and VideoLLaVA [lin2023video], as well as those proposed for driving video analysis DriveMM [huang2024drivemm] and WiseAD [zhang2024wisead]. The dataset and evaluation metric details are provided in the Supplementary Material. All the following evaluations are in a strictly zero-shot setting, with no fine-tuning.
@@
We analyze the performance on accident-cause and traffic scene understanding on MM-AU and SUTD datasets under multi-choice VQA setup, shown in Table˜2 and Table˜2, and gain three insights on $\bm{\dot{\imath}}$Finder versatility and performance. One, $\bm{\dot{\imath}}$Finder outperforms both general-purpose and driving-specialized models on MM-AU and SUTD without any fine-tuning, highlighting its strong out-of-the-box reasoning capabilities. Two, it achieves the best accuracy in all six SUTD categories, from basic understanding to attribution, showing robust generalization across diverse cognitive tasks. Three, the 39-point gap between $\bm{\dot{\imath}}$Finder and DriveMM on MM-AU underscores the limitations of current domain-specific models and the strength of $\bm{\dot{\imath}}$Finder’s architecture in complex, real-world driving scenarios.
@@
VideoChat2 [li2023videochat] 49.56 42.17
Video-LLaVA [lin2023video] 43.63 38.35
@@
Method U F R C I A
Generalist Models
VideoLLaMA2 [damonlpsg2024videollama2] 49.2 39.0 48.5 53.5 35.8 45.2
VideoChat2 [li2023videochat] 42.5 38.1 43.8 49.2 30.4 42.8
Video-LLaVA [lin2023video] 39.7 37.2 35.8 40.5 31.1 36.4
Driving-specialized Methods
DriveMM [huang2024drivemm] 47.6 38.6 40.1 43.2 38.5 37.7
$\bm{\dot{\imath}}$Finder (Ours) 52.2 43.5 50.2 56.8 39.2 49.6
@@
VideoLLaMA2 [damonlpsg2024videollama2] 36.00 4.15 33.45 26.28
VideoChat2 [li2023videochat] 41.20 6.58 36.81 40.98
Video-LLaVA [lin2023video] 21.00 4.26 26.99 31.23
@@
VideoChat2 [li2023videochat] 58.0 0.62 0.57 0.68
VideoLLaMA2 [damonlpsg2024videollama2] 50.0 0.67 0.50 1.00
@@
DriveMM [huang2024drivemm] 49.0 0.32 0.48 0.24
WiseAD [zhang2024wisead] 50.0 0.67 0.50 1.00
@@
Next, we analyzed all methods under an open-ended VQA setup on the LingoQA dataset in Table˜4 and provided the following insights. One, $\bm{\dot{\imath}}$Finder achieves the best Lingo-Judge accuracy among all evaluated methods, outperforming VideoLLaMA2 by 3%, VideoChat2 by 8%, and Video-LLaVA by 23.2%. Further, it outperforms driving-specialized models, WiseAD by 30.8%. Two, while our method performs competitively on BLEU and METEOR metrics, it slightly lags behind VideoChat2. This suggests that VideoChat2 generates responses that are more lexically similar to reference answers but do not necessarily reflect greater factual correctness. We also show this in Figure˜5.
Finally, we analyze all methods for accident occurrence prediction on the Nexar dataset in Table˜4. $\bm{\dot{\imath}}$Finder achieves the highest accuracy at 62.0%, outperforming both generalist models, as well as driving-specialized models. Notably, while VideoLLaMA2 and WiseAD achieve high F1 and perfect recall by always predicting accidents, this limits real-world reliability. In contrast, VideoChat2 balances precision and recall for better accuracy. Building on this, $\bm{\dot{\imath}}$Finder integrates it within the peer V-VLM to enhance both accuracy and precision.